The sound level at a point 5.0 m away from a point source is 40 dB. What will be the level at a point
50 m away from the source?

Grade:12

1 Answers

Arun
6 years ago
We know that β = 10 log base 10 (l/l base 0) Β base A 10 log l base A/l base o, β base B = 10 log l base B/l base o ⇒l base A / l base 0 = 10^( β base B/10) ⇒ l base B/l base o = 10^( β base B/10) ⇒l base A/l base B = r^2 base B/ r^2 base A = (50/5)^2 ⇒ 10^( β base A β base B) = 10^2⇒β base A - β base B/10 = 2 ⇒ β base A - β base b = 20 ⇒β base B = 40 – 20 = 20 dβ.